Which of the following is true about the graph of the function y = 2.5x? Select all that apply
The equation does NOT represent a proportional relationship.
A change of 5 units in x results in a change of 2 units in y.
The unit rate of the relationship is 2/5
The slope of the line is 2.5
A change of 2 units in x results in a change of 5 units in y.

Answer by Theo(13342)

your options are:

1.
The equation does NOT represent a proportional relationship.
this is false since the equation does represent a proportional relationship because y is always 2.5 times as large as x.

2.
A change of 5 units in x results in a change of 2 units in y.
this is false since a change in 5 units of x will result in 2.5 * 5 = a change of 12.5 units in y.

3.
The unit rate of the relationship is 2/5.
since the relationship is the value of y based on the value of x, then the unit rate is 2.5, making this statement false.

4.
The slope of the line is 2.5
this is true, since the slope is the change in y divided by the change in x.
for every change in x of 1 unit, the change in y is 2.5 units.
that makes this statement true.

5.
A change of 2 units in x results in a change of 5 units in y.
this statement is also true, since a change in 2 units of x results in a change of 2 * 2.5 = 5 units in y.

so, it looks like statements 1,2,3 are false and statements 4,5 are true, as best i can determine.

note that the slope intercept form of the equation of a straight line is y = mx + b
m is the slope.
b is the y-intercept.
since the equation is y = 2.5x, then the slope is 2.5 and the y-intercept is 0.